OpenStudy (anonymous):

in terms of triangles.... find a if b=13, and theta =45 degrees. with the equation 1/2absin theta

OpenStudy (cwrw238):

wheres the '=' in your equation?

OpenStudy (anonymous):

there is none.

OpenStudy (cwrw238):

there must be an '=' otherwise its not an equation

OpenStudy (cwrw238):

what you can do is simplify your expression by substituting for b and theta

OpenStudy (anonymous):

thats what I'm given.

OpenStudy (cwrw238):

you omitted to say the area = 17 so 1/2absin theta = 17 (1/2) 13 a sin 45 = 17 a sin 45 = 17 / 6.5 a = 17/ (6.5* sin 45) = 3 . 70

OpenStudy (anonymous):

3.7 didn't work. btw, flaw in your math. 17/6.5 is not 6.5

OpenStudy (anonymous):

maybe thats my bad.

OpenStudy (anonymous):

17/6.5 is 2.62

OpenStudy (cwrw238):

no a = 17 divided by (6.5 * sin 45) not 17 / 6.5 - you work out whats in the parentheses first

OpenStudy (cwrw238):

i'll check it out again..

OpenStudy (cwrw238):

its correct 3.70 correct to 3 decimal places

OpenStudy (cwrw238):

PEDMAS is a useful acronym to remind you of the order of operations in a maths expression it means you calculate in following order parenthesis exponent division/multiply add/subtract so for example 3( 4 + 5)^2 - 2 = 3 * 9^2 - 2 - work out parentheses = 3 * 81 - 2 now exponent 9^2 = 243 - 2 multiplication = 241 subtraction

OpenStudy (cwrw238):

you need any more help?